Kai Stephens recently opened a mobile dog grooming business named Waggin' Clean in the Dallas suburb of Frisco, TX. As a relatively new competitor in the market, Kai knows that he must find more customers quickly if his business is going to survive. There are many competitors in the Dallas market with service prices ranging between $40 and $100. To position his business competitively, Kai has decided to charge $60 for his dog grooming services. Kai's business has fixed monthly expenses of $2000 which covers his salary and the payment on his mobile grooming van. Additionally, Kai has variable expenses of $20 per service for his supplies, variable labor expenses, and gas.

Even with his competitive price, Kai's business is not reaching its potential. Kai has the ability to service 8 canine clients per day, 6 days a week, which results in approximately 192 potential dog services per month. Unfortunately, Kai is only booking 4 grooming appointments on average each day. Kai knows he needs to do some marketing to increase his client base. To increase his bookings, Kai is deciding between two different promotional marketing campaigns.

Kai is deciding between running an advertisement in a local Dallas magazine and creating a $10 off coupon direct mail campaign. The magazine ad Kai is considering has an advertising cost of $1000 while the advertising cost for the direct mailing is $500. Based on his research, Kai believes that the magazine advertisement would result in a 25 percent increase in customers (120 customers per month) while the direct mail coupon would result in a 55 percent increase in customers (150 customers per month).

The goal of this activity is to understand the profit equation and the relationship between sales, revenue, and costs. Keep in mind that:

Profit = Total Revenue - Total Costs. Total Revenue (TR) = Price (P) x Quantity Sold (Q).

1. Based on the data provided in the marketing dashboard, what type of promotional marketing effort, if any, should Kai invest in to increase the number of canine customers and profit?
Answer #1

Answer:

Current data:

Number of canine clients per day = 4

Number of canine clients per month = 4 * 6 * 4 = 96

Service Revenue per canine client = $60

Variable cost = $20 per service

Fixed monthly expenses = $2000

Hence:

Total revenue = 96 * $60 = $5760

Total expenses = Variable cost + fixed cost = 96 * 20 + 2000 = $3920

Monthly Profit = Total Revenue - Total Costs = 5760 - 3920 = $1,840

Now let us evaluate the two promotional marketing options:

Advertisement in a local Dallas magazine:

Number of customers per month =120

Increase in fixed cost per month = $1,000

Hence:

Total revenue = 120 * 60 = $7,200

Total expenses = 120 * 20 + 2000 + 1000 = $5400

Total profit = 7200 - 5400 =$1,800

As such, with advertisement in Dallas magazine, we observe a reduction of profit to $1,800 from current profit (with promotion) of $1840.

Hence this promotion campaign has of be rejected.

Direct mail campaign:

Customer per month increases to = 150

This campaign involves a $10 off coupon and  advertising cost of $500.

With $10 off, Service Revenue per canine client = $60 - 10 = $50

Total revenue = 150 * $50 = $7,500

Total cost = 150 * $20 + 2000 + 500 = $5,500

Profit = 7500 - 5500 = $2,000

Hence:

Hence direct mail promotion campaign result in net incremental profit of (2000 - 1800=) $200

As such:

Kai should invest in direct mail promotion campaign to increase the number of canine customers and profit